d[f(x)]/dx=1/1+x2 => d[f(x)3]/dx=?


a)x square/1+x6


b)x3/1+x6


c)3x2/1+x6


d)3x3/1+x6


Plz explain...

Grade:12

1 Answers

Chetan Mandayam Nayakar
12 years ago

f(x)=arctan(x), d(f(x3))/dx =(1/(1+(x3)2))(3x2), answer= c)